OpenStudy (anonymous):

Please place the following in y=mx+b format. No decimals please also place all fractions as improper and in lowest terms.Also please explan how you got the answer. y-7=1/4(x-3)

OpenStudy (anonymous):

y-7 = 1/4(x-3) y=1/4x- 3/4 + 7 y = 1/4x + 25/4 You can convert 25/4 into an improper fraction

OpenStudy (anonymous):

Thank you soo much